# Multi-modal Gaussian Process Variational Autoencoders for Neural and Behavioral Data

### v1

@@ Abstract
Characterizing the relationship between neural population activity and behavioral data is a central goal of neuroscience. While latent variable models (LVMs) are successful in describing high-dimensional time-series data, they are typically only designed for a single type of data, making it difficult to identify structure shared across different experimental data modalities. Here, we address this shortcoming by proposing an unsupervised LVM which extracts temporally evolving shared and independent latents for distinct, simultaneously recorded experimental modalities. We do this by combining Gaussian Process Factor Analysis (GPFA), an interpretable LVM for neural spiking data with temporally smooth latent space, with Gaussian Process Variational Autoencoders (GP-VAEs), which similarly use a GP prior to characterize correlations in a latent space, but admit rich expressivity due to a deep neural network mapping to observations. We achieve interpretability in our model by partitioning latent variability into components that are either shared between or independent to each modality. We parameterize the latents of our model in the Fourier domain, and show improved latent identification using this approach over standard GP-VAE methods. We validate our model on simulated multi-modal data consisting of Poisson spike counts and MNIST images that scale and rotate smoothly over time. We show that the multi-modal GP-VAE (MM-GPVAE) is able to not only identify the shared and independent latent structure across modalities accurately, but provides good reconstructions of both images and neural rates on held-out trials. Finally, we demonstrate our framework on two real world multi-modal experimental settings: Drosophila whole-brain calcium imaging alongside tracked limb positions, and Manduca sexta spike train measurements from ten wing muscles as the animal tracks a visual stimulus.
@@ 1 Introduction
Recent progress in experimental neuroscience has enabled researchers to record from a large number of neurons while animals perform naturalistic behaviors in sensory-rich environments. An important prerequisite to analyzing these data is to identify how the high-dimensional neural data is related to the corresponding behaviors and environmental settings. Traditionally, researchers often employ a two-step approach, involving first dimensionality reduction on neural data followed by a post-hoc investigation of the latent structure of neural recordings with respect to experimental variables of interest such as stimuli or behavior [1, 2, 3, 4, 5]. Researchers may also specifically design latent variable models (LVMs) that extract low-dimensional structure in neural data using these experimental variables of interest [6, 7, 8, 9]. However, recent deep learning advancements allow for both experimental and behavioral variables to be part of a single latent-variable model [10, 11, 12, 13], thus opening a new avenue for unsupervised discoveries on the relationship between neural activity and behavior.
@@
The existing approaches that jointly model neural activity and behavior are limited in that they often rely on a single latent space to describe data from both modalities [10, 11], making it difficult for practitioners to isolate latent features that are shared across and independent to neural activity or behavior. Approaches that do isolate modality specific and shared latent structure either do so with no temporal structure [14, 15, 13, 16, 17, 18], or use a relatively inflexible linear dynamical system [12]. Moreover, because of the deep neural network mapping from latents to observations, such existing multi-modal approaches generate latent spaces that are not obviously related to experimental variables of interest, and so additional ad hoc model features are often added to aid in interpretability and analysis in neuroscience settings [10, 11, 8].
@@
However, LVMs developed for neural data often are able to uncover interpretable features using neural activity alone in an unsupervised fashion with minimal a priori assumptions. One example of these is Gaussian Process Factor Analysis (GPFA), a widely used LVM in neuroscience that finds smoothly evolving latent dynamics in neural population activity, and can illustrate different aspects of neural processing [19, 20, 21, 22, 23]. GPFA constrains the latent space of neural activity through the use of a Gaussian Process (GP) prior. GP priors have also been adapted to regularize the latent space of variational autoencoders (GP-VAE or GP-prior VAE) with a variety of applications [24, 25, 26, 27]. In each of these approaches, the GP prior provides a flexible constraint in the latent dimension, specifying correlations across auxiliary observations like viewing angle, lighting, or time. The GP prior often is used for out-of-sample prediction in GP-VAEs, but in GPFA is frequently used to visualize latent structure in noisy neural population activity on a trial-by-trial basis [20, 2, 21]. Here, we wish to leverage the interpretability seen in unsupervised GPFA models with the power of GP-VAEs to use with multi-modal time-series datasets in neuroscience.
@@
We propose a model for jointly observed neural and behavioral data that partitions shared and independent latent subspaces across data modalities while flexibly preserving temporal correlations using a GP prior. We call this the multi-modal Gaussian Process variational autoencoder (MM-GPVAE). Our first innovation is to parameterize the latent space time-series in terms of a small number of Fourier frequencies, an approach that has been used before in the linear GPFA setting [28, 29, 21]. We show that the Fourier representation dramatically improves latent identification for the standard GP-VAE. Our second innovation is to augment our new Fourier GP-VAE model to describe two data modalities simultaneously by mapping the latents linearly to neural data (like in GPFA), as well as nonlinearly to another experimental variable via a deep neural network (like a GP-VAE). We leave the specific identity of this other experimental modality intentionally vague - it could be keypoints of limb positions as animals freely move or visual stimuli across time. Because this observation modality is characterized by a deep neural network, our model can be adapted to any additional experimental variable of interest.
@@
We validate our MM-GPVAE model on a simulated dataset of a smoothly rotating and scaling MNIST digit alongside simulated Poisson neural activity, whereby the digit data and the spiking data share latent structure. We show that the MM-GPVAE model is able to recover both shared and independent latent structures while simultaneously providing accurate reconstructions of data from both modalities. Lastly, we demonstrate the utility of our model by fitting the MM-GPVAE to two real-world multi-modal neural datasets: 1) Drosophila (fly) whole-brain calcium imaging alongside tracked 16 limb positions, and 2) Manduca sexta (hawkmoth) spike train measurements from ten wing muscles alongside a continuously moving visual stimulus. In the former case, we demonstrate that the neural and shared subspaces best separate Drosophila behavioral conditions and in the latter case, we show distinct time-varying latent components tracking muscle and stimuli movement in the experiment. By showing the MM-GPVAE in these two domains, we demonstrate that our model is adaptable to a range of diverse experimental preparations in systems neuroscience.
@@ 2 The Gaussian Process Variational Autoencoder
The Gaussian Process variational autoencoder (GPVAE) uses high-dimensional data (e.g. images) accompanied by auxiliary information, like viewing angle, lighting, object identity, or observation time. This auxiliary information provides the indices in the GP prior latent representation, specifying correlations across the latent space and allowing for out-of-sample predictions at new auxiliary values [24, 25, 26]. However, in our case we consider continuously observed time-series data in experimental neuroscience experiments, so our auxiliary information here are evenly sampled time-bins. Because of this, we can leverage the advantages of the Fourier-domain GP representation [29, 22, 28].
Formally, consider smoothly-varying image data across timepoints, represented by the pixels-by-time matrix . For latent variable each latent () evolves according to a Gaussian process, , with covariance kernel . The time-by-time covariance matrix of each is then given by the Gram matrix corresponding to . We use a squared exponential (RBF) kernel for governed by a marginal variance and length scale with an additive diagonal term to help with inference [30, 24]. The likelihood of the image data is Gaussian with mean given by the latent values at any timepoint passed through a deep neural network with parameters and whose covariance is .
The GP-VAE is learned using standard VAE amortized inference [31], where the parameters and of a variational distribution are given as neural network functions of the observed data . Here, the evidence lower bound (ELBO) is maximized with respect to the variational parameters , model parameters , and GP hyperparameters. In this work, is set to a fixed value of for all experiments except for the final data analysis example, where it is set to a value of . While the ELBO may be expressed in a variety of ways, we will follow the form that includes the variational entropy term. For details about this approach for the standard GP-VAE, see [24].
@@ 2.1 Fourier-domain representation of the GP-VAE
We consider a version of the GP-VAE whose auxiliary variables are a Fourier frequency representation of the time domain, as opposed to timepoints sampled on a regular lattice. This allows us to parameterize a frequency representation of the latent variables that is Gaussian distributed according to a where the original RBF GP covariance matrix may be diagonalized by Here, is the orthonormal discrete Fourier transform matrix and represents the frequency. The model prior can now be written as
Where represents a -values long vector of Fourier frequencies, and represents a matrix of the frequency representation of the P latent variables at the frequencies. The model likelihood retains the same form as the standard GP-VAE.
@@
where indexes a trial of times-series images . Training is done by batches over a subset of trials. In contrast to the standard GP-VAE [24], the Fourier representation requires that, for each trial, images at all timepoints are mapped to a single -dimensional Fourier representation. We accomplish this in two steps - first, we use a deep neural network for each image at each time point , which will result in total network embeddings for a single trial, each of dimension . We follow that with a single linear layer each for the mean and variance of the variational distribution, , , that will map from the number of timepoints to the number of Fourier frequencies. . A schematic for the Fourier domain variational and generative architecture is displayed in Figure 1(a).
There are a number of advantages to representing the GP-VAE latents and variational parameters in the Fourier domain. For one, the diagonal representation of means we avoid a costly matrix inversion when evaluating the GP prior [29, 28, 32, 33, 34]. Secondly, we can prune the high frequencies in the Fourier domain, effectively sparsifying the variational parameters, and hence frequencies (), while enforcing a smooth latent representation. Though this is a free parameter in our model thus if the user would like to model steady-states they are able to easily adjust this. Lastly, the Fourier representation of the variational parameters mean we can retain both temporal correlations and the advantages of using a mean-field approximation for the variational posterior [21].These advantages have been seen in simpler GP models. For more information, see [29, 21].
The Fourier-represented GP-VAE dramatically improves the ability of the GP-VAE to learn a true underlying smoothly evolving latent for high-dimension data in a non-linear model. We demonstrate this on a simulated example using an MNIST digit that rotates by a time-varying angle given by a draw from a GP with an RBF kernel. We fit our Fourier GP-VAE model as well as the standard VAE and standard GP-VAE models to these trials. In each case, we are looking to recover the true underlying generative latent angle and hence we assert a one-dimensional latent dimensionality for the model. For more information on training and testing, see the supplement.
Because the Fourier variational distribution preserves temporal correlations and prunes the high-frequency components of the latents , the Fourier GP-VAE model shows much smoother underlying latent representations on held-out trials (Fig 1 (c)), though each model retains the ability to accurately reconstruct the images through the network mapping from the latent space (Fig 1(b)). When the latent space is mapped through an affine transform to align latents on held-out trials to the true latent space, the Fourier domain GP-VAE does much better uncovering the true smoothly evolving latent angle (Fig 1 (c), (d)).
@@ 3 The Multi-Modal Gaussian Process Variational Autoencoder
We now focus on extending the GP-VAE to model data of two modalities simultaneously. The examples we will emphasize here involve neural activity alongside some other experimental variable such as naturalistic movement or high-dimensional stimuli. The observations of our model are two distinct data modalities represented by matrices and . Here, denotes time-bin indices, denotes trials, and and denote the dimension of the observations for each modality. We assume, as before, that all data are generated by smoothly varying low-dimensional latent variables that are now either independent to, or shared between data modalities. As before, latents are initially parameterized in a pruned Fourier representation before being mapped to the time domain. The latents are then partitioned by a loadings matrix which linearly combines the shared latent representation with the independent latents.
Here, refers to latents that are shared between regions, while and denote region-specific latent variables and represents an additive offset. The outputs after this linear mixing result in modality-specific embeddings which we call and . Here, modality , the ’behavioral modality’ is passed through a deep neural network, and modality , the ’neural modality’, is passed through a pointwise nonlinearity to enforce non-negativity of Poisson rates. The likelihood of the data given the embedding is:
Where the function represents a decoder neural network for the experimental data and we use the exponential nonlinearity for . Learning is performed as before - the mean and variance of a mean-field Gaussian variational distribution is given by the data passed through a neural network with parameters . Here, our lower bound is similar to that found in Casale et al. (2018) with an additional term for the neural data modality.
The expectation with respect to the variational distribution is performed by drawing stochastic samples from a standard normal and using the variational mean and variance to get a sample from (the so-called reparameterization trick [31]). However, because we use a mean-field Gaussian variational distribution, both the Poisson log-likelihood and the GP prior terms in the expectation can be computed in closed-form and preclude the need for sampling for these terms. For clarity, we omit that closed-form expression here. See the supplment for details.
@@ 4.1 Simulated data
We first assess the performance of the MM-GPVAE model using a simulated dataset of two modalities: a smoothly rotating and scaling MNIST digit, and simultaneously accompanied Poisson spike counts from 100 neurons. A total of three latents were drawn from an underlying GP kernel to generate single-dimensional shared, neural, and image subspaces. There were a total of 300 trials, each trial consisting of 60 timebins. The data was split into 80% for training and 20% for testing. For this simulated example, one latent represents an interpretable modulation of the image as it directly effects the scaling of the MNIST digit. So as the latent values change along the trial, the image scales smaller and larger continuously. Another latent, corresponding to the independent component of the neural modality, provides one component of the log-rates of the Poisson spiking data. The final latent reflects the shared variability between both modalities. Here, the latent is again interpretable with respect to the image in that this latent corresponds to the angle of the smoothly rotating MNIST digit. This shared latent also linearly combines with the neural-only subspace to provide the the log-rates of the spiking Poisson population.
@@
The MM-GPVAE trained simultaneously on both images and spikes can successfully recover the three-dimensional latent structure across these two modalities. Figure 3(a) shows the underlying true latent in a held-out trial alongside the estimated latent structure extracted from the MM-GPVAE. As before, the held-out latent space is scaled to align with the true latent space as closely as possible, as the MM-GPVAE model latent space is invariant to scaling transform. In addition to accurately extracting the latent structure, the MM-GPVAE also has the ability to accurately reconstruct both neural rates and images across the time-series from the latent space. Four example images and three example neural rates are show in Figure 3 (b). For more examples, see the supplemental materials.
@@
The MM-GPVAE is also able to exploit information across modalities to better model the observed data. The left side of 3(c) shows the mean squared error (MSE) across pixels for reconstructed images in held-out trials. By leveraging latent information from the spiking modality, the MM-GPVAE is able to reconstruct the images better than a single-modality GP-VAE trained on the images alone [24]. Similarly, the held-out neural rates are better estimated when the image data is included, though the effect here is more modest than for the image modality. Finally, we also show that the shared latent variable can be more accurately identified when data from both modalities is used (Figure 3(c)). Here we assess the accuracy of the shared latent estimate on test trials when we use data from each modality alone (GPVAE and GPFA) compared to both modalities simultaneously (MM-GPVAE). As expected, data from both modalities better identifies shared structure across the data types. The settings for all three models were exactly the same for these comparisons. We only ablated the models to make it into uni-modal settings. For additional details regarding training and the neural network set-ups for this example, please see the supplemental material.

### v2

@@ Abstract
Characterizing the relationship between neural population activity and behavioral data is a central goal of neuroscience. While latent variable models (LVMs) are successful in describing high-dimensional time-series data, they are typically only designed for a single type of data, making it difficult to identify structure shared across different experimental data modalities. Here, we address this shortcoming by proposing an unsupervised LVM which extracts temporally evolving shared and independent latents for distinct, simultaneously recorded experimental modalities. We do this by combining Gaussian Process Factor Analysis (GPFA), an interpretable LVM for neural spiking data with temporally smooth latent space, with Gaussian Process Variational Autoencoders (GP-VAEs), which similarly use a GP prior to characterize correlations in a latent space, but admit rich expressivity due to a deep neural network mapping to observations. We achieve interpretability in our model by partitioning latent variability into components that are either shared between or independent to each modality. We parameterize the latents of our model in the Fourier domain, and show improved latent identification using this approach over standard GP-VAE methods. We validate our model on simulated multi-modal data consisting of Poisson spike counts and MNIST images that scale and rotate smoothly over time. We show that the multi-modal GP-VAE (MM-GPVAE) is able to not only identify the shared and independent latent structure across modalities accurately, but provides good reconstructions of both images and neural rates on held-out trials. Finally, we demonstrate our framework on two real-world multi-modal experimental settings: Drosophila whole-brain calcium imaging alongside tracked limb positions, and Manduca sexta spike train measurements from ten wing muscles as the animal tracks a visual stimulus.
@@ 1 Introduction
Recent progress in experimental neuroscience has enabled researchers to record from a large number of neurons while animals perform naturalistic behaviors in sensory-rich environments. An important prerequisite to analyzing these data is to identify how the high-dimensional neural data is related to the corresponding behaviors and environmental settings. Traditionally, researchers often employ a two-step approach, involving first dimensionality reduction on neural data followed by a post-hoc investigation of the latent structure of neural recordings with respect to experimental variables of interest such as stimuli or behavior (Wu et al., 2017; Zhao and Park, 2017; Saxena and Cunningham, 2019; Bahg et al., 2020; Pei et al., 2021). However, recent deep learning advancements allow for both experimental and behavioral variables to be part of a single latent-variable model (Schneider et al., 2023; Hurwitz et al., 2021; Sani et al., 2021; Singh Alvarado et al., 2021), thus opening a new avenue for unsupervised discoveries on the relationship between neural activity and behavior.
@@
The existing approaches that jointly model neural activity and behavior are limited in that they often rely on a single latent space to describe data from both modalities (Schneider et al., 2023; Hurwitz et al., 2021), making it difficult for practitioners to isolate latent features that are shared across and independent to neural activity or behavior. Approaches that do isolate modality specific and shared latent structure either do so with no temporal structure (Gundersen et al., 2019; Kleinman et al., 2024; Singh Alvarado et al., 2021; Wu and Goodman, 2018; Shi et al., 2019; Brenner et al., 2022), or use a relatively inflexible linear dynamical system (Sani et al., 2021). Moreover, because of the deep neural network mapping from latents to observations, many existing multi-modal approaches generate latent spaces that are not obviously related to experimental variables of interest, and so additional model features are often added to aid in interpretability and analysis in neuroscience settings (Schneider et al., 2023; Hurwitz et al., 2021; Zhou and Wei, 2020).
@@
However, LVMs developed for neural data often are able to uncover interpretable features using neural activity alone in an unsupervised fashion with minimal a priori assumptions. One example of these is Gaussian Process Factor Analysis (GPFA), a widely used LVM in neuroscience that finds smoothly evolving latent dynamics in neural population activity, and can illustrate different aspects of neural processing (Yu et al., 2008; Duncker and Sahani, 2018; Keeley et al., 2020a; b; Balzani et al., 2022). GPFA constrains the latent space of neural activity through the use of a Gaussian Process (GP) prior. GP priors have also been adapted to regularize the latent space of variational autoencoders (GP-VAE or GP-prior VAE) with a variety of applications (Casale et al., 2018; Fortuin et al., 2020; Ramchandran et al., 2021; Yu et al., 2022). In each of these approaches, the GP prior provides a flexible constraint in the latent dimension, specifying correlations across auxiliary observations like viewing angle, lighting, or time. The GP prior often is used for out-of-sample prediction in GP-VAEs, but in GPFA is frequently used to visualize latent structure in noisy neural population activity on a trial-by-trial basis (Duncker and Sahani, 2018; Zhao and Park, 2017; Keeley et al., 2020a). Here, we wish to leverage the interpretability seen in unsupervised GPFA models with the power of GP-VAEs to use with multi-modal time-series datasets in neuroscience.
@@
We propose a model for jointly observed neural and behavioral data that partitions shared and independent latent subspaces across data modalities while flexibly preserving temporal correlations using a GP prior. We call this the multi-modal Gaussian Process variational autoencoder (MM-GPVAE). Our first innovation is to parameterize the latent space time-series in terms of a small number of Fourier frequencies, an approach that has been used before in the linear GPFA setting (Paciorek, 2007; Aoi and Pillow, 2017; Keeley et al., 2020a). We show that the Fourier representation dramatically improves latent identification for the standard GP-VAE. Our second innovation is to augment our new Fourier GP-VAE model to describe two data modalities simultaneously by mapping the latents linearly to neural data (like in GPFA), as well as nonlinearly to another experimental variable via a deep neural network (like a GP-VAE). We leave the specific identity of this other experimental modality intentionally vague - it could be keypoints of limb positions as animals freely move or visual stimuli across time. Because this observation modality is characterized by a deep neural network, our model can be adapted to any experimental variable of interest.
@@
We validate our MM-GPVAE model on a simulated dataset of a smoothly rotating and scaling MNIST digit alongside simulated Poisson neural activity, whereby the digit data and the spiking data share latent structure. We show that the MM-GPVAE model is able to recover both shared and independent latent structures while simultaneously providing accurate reconstructions of data from both modalities. Lastly, we demonstrate the utility of our model by fitting the MM-GPVAE to two real-world multi-modal neural datasets: 1) Drosophila (fly) whole-brain calcium imaging alongside tracked 16 limb positions, and 2) Manduca sexta (hawkmoth) spike train measurements from ten wing muscles alongside a continuously moving visual stimulus. In the former case, we are able to view Drosophila behavioral conditions across shared and independent subspaces and in the latter case, we show distinct time-varying latent components tracking muscle and stimuli movement in the experiment. By showing the MM-GPVAE in these two domains, we demonstrate that our model is adaptable to a range of diverse experimental preparations in systems neuroscience.
@@ 2 The Gaussian Process Variational Autoencoder
The Gaussian Process variational autoencoder (GPVAE) uses high-dimensional data (e.g. images) accompanied by auxiliary information, like viewing angle, lighting, object identity, or observation time. This auxiliary information provides the indices in the GP prior latent representation, specifying correlations across the latent space and allowing for out-of-sample predictions at new auxiliary values (Casale et al., 2018; Fortuin et al., 2020; Ramchandran et al., 2021). However, in our case we consider continuously observed time-series data in experimental neuroscience experiments, so our auxiliary information here is evenly sampled time-bins. Because of this, we can leverage the advantages of the Fourier-domain GP representation (Aoi and Pillow, 2017; Keeley et al., 2020b; Paciorek, 2007).
Formally, consider smoothly-varying image data across timepoints, represented by the pixels-by-time matrix . For latent variable each latent () evolves according to a Gaussian process, , with covariance kernel . The time-by-time covariance matrix of each is then given by the Gram matrix corresponding to . We use a squared exponential (RBF) kernel for governed by a marginal variance and length scale with an additive diagonal term to help with inference (Yu et al., 2009; Casale et al., 2018). The likelihood of the image data is Gaussian with mean given by the latent values at any timepoint passed through a deep neural network with parameters and whose covariance is .
The GP-VAE is learned using standard VAE amortized inference (Kingma and Welling, 2013), where the parameters and of a variational distribution are given as neural network functions of the observed data , parameterized by . Here, the evidence lower bound (ELBO) is maximized with respect to the variational parameters , model parameters , and GP hyperparameters. In this work, is set to a fixed value of for all experiments except for the final data analysis example, where it is set to a value of . While the ELBO may be expressed in a variety of ways, we will follow the form that includes the variational entropy term. For details about this approach for the standard GP-VAE, see Casale et al. (2018).
@@ 2.1 Fourier-domain representation of the GP-VAE
We consider a version of the GP-VAE whose auxiliary variables are a Fourier frequency representation of the time domain, as opposed to timepoints sampled on a regular lattice. This allows us to parameterize a frequency representation of the latent variables that is Gaussian distributed according to where the original RBF GP covariance matrix may be diagonalized by Here, is the orthonormal discrete Fourier transform matrix and is a single Fourier-domain latent vector whose length is equal to , the total number of Fourier frequencies representing the latent. The model prior can now be written as
Where represents a matrix of the frequency representation of the latent variables at the frequencies and is an covariance matrix. The model likelihood retains the same form as the standard GP-VAE,
@@
where indexes a trial of times-series images and indexes the Fourier frequencies (). In contrast to the standard GP-VAE (Casale et al., 2018), the Fourier representation requires that, for each trial, images at all timepoints are mapped to a single -dimensional Fourier representation. We accomplish this in two steps - first, we use a deep neural network for each image at each time point , which will result in total network embeddings for a single trial, each of dimension . We follow that with a single linear layer each for the mean and variance of the variational distribution, , , that will map from the number of timepoints to the number of Fourier frequencies. . A schematic for the Fourier domain variational and generative architecture is displayed in Figure 1(a).
There are a number of advantages to representing the GP-VAE latents and variational parameters in the Fourier domain. For one, the diagonal representation of means we avoid a costly matrix inversion when evaluating the GP prior (Aoi and Pillow, 2017; Paciorek, 2007; Hensman et al., 2017; Wilson and Adams, 2013; Loper et al., 2021). Secondly, we can prune the high frequencies in the Fourier domain, effectively sparsifying the variational parameters, and hence frequencies (), while enforcing a smooth latent representation. Lastly, the Fourier representation of the variational parameters mean we can retain both temporal correlations as well as the advantages of using a mean-field approximation for the variational posterior .These advantages have been seen in simpler GP models. For more information, see (Aoi and Pillow, 2017; Keeley et al., 2020a).
The Fourier-represented GP-VAE dramatically improves the ability of the GP-VAE to learn a true underlying smoothly evolving latent for high-dimension data in a non-linear model. We demonstrate this on a simulated example using an MNIST digit that rotates by a time-varying angle given by a draw from a GP with an RBF kernel. We fit our Fourier GP-VAE model as well as the standard VAE and standard GP-VAE models to these trials. In each case, we are looking to recover the true underlying generative latent angle and hence we assert a one-dimensional latent dimensionality for the model. For more information on training and testing, see appendix.
Because the Fourier variational distribution preserves temporal correlations and prunes the high-frequency components of the latents , the Fourier GP-VAE model shows much smoother underlying latent representations on held-out trials (Fig 1 (c)), though each model retains the ability to accurately reconstruct the images through the network mapping from the latent space (Fig 1(b)). When the latent space is mapped through an affine transform to align latents on held-out trials to the true latent space, the Fourier domain GP-VAE does much better uncovering the true generative latent angle (Fig 1 (c), (d)).
@@ 3 The Multi-Modal Gaussian Process Variational Autoencoder
We now focus on extending the GP-VAE to model data of two modalities simultaneously. The examples we will emphasize here involve neural activity alongside some other experimental variable such as naturalistic movement or high-dimensional stimuli. The observations of our model are two distinct data modalities represented by matrices and . Here, denotes time-bin indices, denotes trials, and and denote the dimension of the observations for each modality. We assume, as before, that all data are generated by smoothly varying low-dimensional latent variables that are now either independent to or shared between data modalities. As before, latents are initially parameterized in a pruned Fourier representation before being mapped to the time domain. The latents are then partitioned by a loadings matrix which linearly combines the shared latent representation with the independent latents.
Here, refers to latents that are shared between modalities, while and denote modality-specific latent variables and represents an additive offset. The outputs after this linear mixing result in modality-specific embeddings which we call and . Here, modality , the "neural modality", is passed through a pointwise nonlinearity to enforce non-negativity of Poisson rates, and modality , the "behavioral modality", is passed through a deep neural network. The likelihood of the data given the embedding is:
Where the function represents a decoder neural network for the behavioral data and we use the exponential nonlinearity for . Learning is performed as before - the mean and variance of a mean-field Gaussian variational distribution is given by the data passed through a neural network with parameters . Here, our lower bound is similar to that found in Casale et al. (2018) with an additional term for the neural data modality.
| ELBO | (6) | |||
@@ 4.1 Simulated data
We first assess the performance of the MM-GPVAE model using a simulated dataset of two modalities: a smoothly rotating and scaling MNIST digit, and simultaneously accompanied Poisson spike counts from 100 neurons. A total of three latents were drawn from an underlying GP kernel to generate single-dimensional shared, neural, and image subspaces. There were a total of 300 trials, each trial consisting of 60 timebins. The data was split into 80% for training and 20% for testing. For this simulated example, one latent represents an interpretable modulation of the image as it directly affects the scaling of the MNIST digit. So as the latent values change along the trial the image scales smaller and larger continuously. Another latent, corresponding to the independent component of the neural modality, provides one component of the log rates of the Poisson spiking data. The final latent reflects the shared variability between both modalities. Here, the latent is again interpretable with respect to the image in that this latent corresponds to the angle of the smoothly rotating MNIST digit. This shared latent also linearly combines with the neural-only subspace to provide the log rates of the spiking Poisson population.
@@
The MM-GPVAE trained simultaneously on both images and spikes can successfully recover the three-dimensional latent structure across these two modalities. Figure 3(a) shows the underlying true latent in a held-out trial alongside the estimated latent structure extracted from the MM-GPVAE. As before, the held-out latent space is scaled to align with the true latent space as closely as possible, as the MM-GPVAE model latent space is invariant to scaling transform. In addition to accurately extracting the latent structure, the MM-GPVAE also has the ability to accurately reconstruct both neural rates and images across the time series from the latent space. Four example images and three example neural rates are shown in Figure 3(b). For more examples, see Figure 8 and 9.
@@
The MM-GPVAE is also able to exploit information across modalities to better model the observed data. The left side of Figure 3(c) shows the mean squared error (MSE) across pixels for reconstructed images in held-out trials. By leveraging latent information from the spiking modality, the MM-GPVAE is able to reconstruct the images better than a single-modality GP-VAE trained on the images alone (Casale et al., 2018). This improvement is additionally dependent on the number of neurons analyzed (Fig 12). Similarly, the held-out neural rates are better estimated when the image data is included, though the effect here is more modest. Finally, we also show that the shared latent variable can be more accurately identified when data from both modalities is used (Fig 3(c)). Here we assess the accuracy of the shared latent estimate on test trials when we use data from each modality alone (akin to standard GP-VAE and GPFA) compared to both modalities simultaneously (MM-GPVAE). As expected, data from both modalities better identifies shared structure across the data types. This additionally does not depend on the structure of nature of the encoding distributions, (Fig 10), and is markedly improved by representing the MM-GPVAE with pruned Fourier frequencies (Fig 11).